Problem 18-11 On July 1, 2017, Sheffield Construction Company Inc. contracted to build an office building for Gumbel Corp. fo2018 2019(b) Using the completed-contract method, prepare schedules to compute the profit or loss to be recognized as a result of this
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Sheffield Construction Company
Percentage of completion method
Calculation of Gross Profit
Particulars 2017 2018 2019
Contract Price           1,920,000          1,920,000          1,920,000 A
Cost Incurred to date              273,600          1,171,600          2,127,600 B
Cost Incurred during the year              273,600             898,000             956,000 C
(Cost incurred to date of the present year less Cost incurred to date of the previous year)
Estimated cost to complete           1,246,400               84,840                        -   D
Estimated total cost           1,520,000          1,256,440          2,127,600 E=C+B
(Cost Incurred to date+ Estimated cost to complete )
Percentage complete 18.00% 93.25% 100.00% F=B/E
(Cost Incurred to date/Estimated total cost)
Revenue to be recognized              345,600          1,444,754             129,646 G
(Contract Price*Percentage complete)- revenue previously recognized
Year 2017: 18% completed. Revenue recognized = 18% x $ 1,920,000 = $ 345,600.
Year 2018: 93.25% completed. Revenue recognized = 93.25% x $ 1,920,000 – $ 345,600 (previously recognized) = $ 1,444,754.
Year 2019: 100% completed. Revenue recognized = 100% x $ 1,920,000 – $ 345,600- $ 1,444,754 (previously recognized) = $ 129,646.
Gross Profit/(Loss)                72,000             546,754            (826,354) H=G-C
Completed contract method
Calculation of Gross Profit
Particulars 2019
Contract Price          1,920,000 A
Cost Incurred to date          2,127,600 B
Gross Profit/(Loss)            (207,600) I=A-B